The capacity of an RD bottle is 100 cc and its mass is 45g. Find the mass of the bottle when it is filled with a liquid of density 600 kg/m^3.
1
Expert's answer
2018-09-20T10:21:09-0400
m = ρ * V, where m – weight, ρ – density, V – volume
ρ = 600 kg/m^3 = 0.6 g/cm^3
We can find a mass of liquid when bottle is filled:
m’ = 0.6*100 = 60 g
So, total mass of bottle is:
m = 60 g + 45g = 105 g